OpenStudy (anonymous):

How to verify csc^2theda/cot^2theda-1=sec2theda

OpenStudy (idku):

csc^2theda/cot^2theda = (1/sin x ) ÷ (cos² x / sin² x) = (1/sin² x ) × (sin² x / cos² x) = 1 / cos² x so the left side equals 1/cos² x or same as sec² x So the 2 sides are equal.